OpenStudy (anonymous):

What is the solution to the system of equations? x= 3y+5 x-4y = 3

OpenStudy (anonymous):

(___,___)

OpenStudy (anonymous):

i am confused on these kinds but i will try it out one sec

OpenStudy (anonymous):

x=3? I really have no clue. Hold on

OpenStudy (anonymous):

um ok that is not very helpful plz don't guess

Vocaloid (vocaloid):

start by rewriting the second equation x= 3y+5 x = 4y + 3 then set them equal to each other

Vocaloid (vocaloid):

3y + 5 = 4y + 3 y = ?

OpenStudy (anonymous):

Wait. Is it x=8y and the bottom one for x=7

OpenStudy (anonymous):

hold on one sec let me figure it out

OpenStudy (anonymous):

8:7

OpenStudy (anonymous):

oh y = 2!!!

OpenStudy (anonymous):

Oh yeah!

Vocaloid (vocaloid):

right. now we know that y = 2 x = 3y + 5 = ?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

11

OpenStudy (anonymous):

is the answer 11?

OpenStudy (anonymous):

cause if y=2 wouldn't it be x= 3 times 2+5=11 which is x!!!

OpenStudy (anonymous):

is that it? i was having trouble with the problem

OpenStudy (anonymous):

They both equal 11 by the way

OpenStudy (anonymous):

yeah!!! They both equal 11

OpenStudy (anonymous):

x=11 and y=2 so y represents it going to the 2nd power so you would multiply

OpenStudy (anonymous):

The answer for both is 11

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

Yeah. It is, I just checked

Vocaloid (vocaloid):

the solution is (2,11) since x= 2 and y = 11

OpenStudy (anonymous):

exactly. See, it's simple

OpenStudy (anonymous):

but 1st can u guys help me figure out something

OpenStudy (anonymous):

x= 27 + 347 - 42 divided by 27.6 x=37 divided by 3 +7

OpenStudy (anonymous):

I don't get it at all